How many moles of P4 would you need to produce 34 grams of PH3?
P4 (s) + 6H2 (g) ____> 4PH3 (g)
1
Expert's answer
2013-05-22T11:51:46-0400
First of all we can find the amount of substance of PH3. We can use the formula
n = m/M
where
n - the amount of substance
m - the mass
M - the molar mass, which can be calculated as follows: M(PH3) = Ar(P) +3*(H) = 31+ 3*1 = 34 g/mol
n(PH3) = 34 g / 34 g/mol = 1 moles
According the equation
P4(s) + 6H2(g) ---> 4PH3(g)
we can find the amount of substance of P4:
n(P4) = n(PH3)/4 = 1/4 = 0.25 moles
Answer: 0.25 moles